OpenStudy (anonymous):

A coin is flipped, a die is rolled, and a card is drawn from a standard deck of cards. What is the probability of the following event? P(tails and 3, and king)?

hartnn (hartnn):

can u find the probability of getting a tail when a coin is flipped ? p(tail) = ?

OpenStudy (anonymous):

I'll try, Hold on.

OpenStudy (anonymous):

1/2? idk :/

hartnn (hartnn):

yup, its 1/2. now do u know the probability of getting a 3 when a dice is rolled ?

OpenStudy (anonymous):

I think, Isn't it P(even)=3/6 and P(odd)=1/2?

hartnn (hartnn):

yes, but we want P(3) now since a dice has 1 to 6 faces, and 3 appears only once, the probability of getting a '3' will be 1/6 got this ?

OpenStudy (anonymous):

Got it.

hartnn (hartnn):

so what did u get your final answer as?

OpenStudy (anonymous):

But it also says a king?

OpenStudy (anonymous):

If you mulitply 1/2 and 1/6 you get 1/12

hartnn (hartnn):

yes, king is remaining, there are 4 kings in deck of 52 cards. so probability is 4/52 = 1/13 overall probability = 1/2 * 1/6 * 1/13

OpenStudy (anonymous):

1/156?

hartnn (hartnn):

yup, thats correct

OpenStudy (anonymous):

Yay, Okay. Thanks Again cx

hartnn (hartnn):

welcome again :)